# The Unreasonable Effectiveness of Easy Training Data for Hard Tasks

### v1

@@ Abstract
How can we train models to perform well on hard test data when hard training data is by definition difficult to label correctly? This question has been termed the scalable oversight problem and has drawn increasing attention as language models have continually improved. In this paper, we present the surprising conclusion that current language models often generalize relatively well from easy to hard data, even performing as well as “oracle” models trained on hard data. We demonstrate this kind of easy-to-hard generalization using simple training methods like in-context learning, linear classifier heads, and QLoRA for seven different measures of datapoint hardness, including six empirically diverse human hardness measures (like grade level) and one model-based measure (loss-based). Furthermore, we show that even if one cares most about model performance on hard data, it can be better to collect and train on easy data rather than hard data, since hard data is generally noisier and costlier to collect. Our experiments use open models up to 70b in size and four publicly available question-answering datasets with questions ranging in difficulty from 3rd grade science questions to college level STEM questions and general-knowledge trivia. We conclude that easy-to-hard generalization in LMs is surprisingly strong for the tasks studied, suggesting the scalable oversight problem may be easier than previously thought. [^1] Our code is publicly available at: https://github.com/allenai/easy-to-hard-generalization.
@@ 1 Introduction
Figure 1: A model trained on easy data (e.g., 3rd Grade problems) does almost as well on hard tasks (College problems) as a model trained on those hard tasks (the College bar). Results shown for Mixtral-8x7B prompted with $k=10$ examples, averaged over 5 random seeds.
It is difficult to supervise LMs (i.e., train LMs to give correct outputs) in specialized domains of human knowledge, because it is difficult to correctly label data in such domains. Labeling difficulty manifests itself in both time to annotate (and thus cost) and label noise Lease (2011); Northcutt et al. (2021). Labeling difficulty becomes severe when specific expertise is required Sambasivan et al. (2021). For example, for sufficiently specific physics problems, PhD holders and PhD students can make errors on as many as 40% of (objective) problems Rein et al. (2023). As more NLP benchmarks focus on challenging domain-specific tasks, having access to large human-labeled training corpora for these tasks may become increasingly infeasible (e.g., existing benchmarks like MMLU Hendrycks et al. (2020) and GPQA Rein et al. (2023) do not come with training data). The question arises: how can we train models to solve hard problems when correctly labeling enough hard data for training is difficult? This problem is an example of the scalable oversight problem, which concerns how to give a good reward signal to a model when it is difficult to assess if its outputs are correct Amodei et al. (2016).
In this paper, we study the problem of easy-to-hard generalization, which is relevant for determining how challenging the scalable oversight problem is in the first place. Easy-to-hard generalization refers to model performance on hard test data when trained only on easy training data, defined according to some human hardness measure (like grade level). Since gathering training data in domains like graduate level STEM fields is expensive and time-consuming, it would clearly be useful if we could improve model performance in these domains by only training models on cleanly labeled data from simpler domains, like high school STEM questions. To assess how well current LMs generalize from easy to hard data, we train [^2] By “training” models, we refer to fitting models to data via in-context learning (ICL), parameter efficient finetuning (QLoRA), or by training a linear classifier head. models on easy data and test them on hard data (“easy-to-hard”), then compare them to an “oracle” upper bound and unsupervised lower bound. The “oracle” upper bound is a model that has access to labeled hard data for training (“hard-to-hard”), while the unsupervised lower bound is a model that is prompted zero-shot to answer questions (“unsupervised-to-hard”). The final metric we are interested in is the Supervision Gap Recovered (SGR):
@@
where Easy, Hard, and Unsupervised refer to model performance on hard test data when trained on easy data, hard data, or no data (zero-shot), respectively. This metric takes a value of 100% when training on easy data is as effective as training on hard data, and it is 0% when a model trained on easy data is no better than prompting a model zero-shot.
Figure 2: Supervising Llama-2-70b with easy data (left: high school level computer science problems; right: math problems with 2-3 reasoning steps) can enable generalization to hard data (left: a college level computer science problem; right: a math problem involving 8 reasoning steps).
Our main result is that models generalize surprisingly well from easy to hard data, often performing almost as well as a model trained on hard data (illustrated in Fig. 1). In experiments with ARC Clark et al. (2018), MMLU Hendrycks et al. (2020), GSM8k Cobbe et al. (2021), and StrategyQA Geva et al. (2021), we find that the Supervision Gap Recovered is usually between 70% and 100%, meaning that easy supervision is at least 70% as good as hard supervision for hard test performance (see Fig. 2 for example problems). These results are robust across (1) model scale between 7b and 70b parameters, (2) six different human hardness measures and a model-based measure, (3) four datasets/tasks, and (4) several training methods including in-context learning with and without chain-of-thought reasoning Brown et al. (2020); Wei et al. (2022), QLoRA Dettmers et al. (2023), and linear classifier heads Belinkov (2022). Overall, our results suggest that current LMs generalize relatively well to held-out data across human difficulty levels even when trained on data that is measurably easier than the test data.
This finding implies that easy supervision may outperform hard supervision when (1) within some data collection budget, a greater quantity of easy data can be collected than hard data, or (2) easy data can be labeled with lower error rates than hard data (see Sec. 5.3). In these settings, we are able to train models that perform better on hard test data when trained on easy data rather than hard data.
Lastly, we aim to understand how easy-to-hard generalization changes with model scale and the gap between train and test hardness (Sec. 5.4). We find that (1) the Supervision Gap Recovered is highly robust across model scales between 7b and 70b parameters, and (2) easy-to-hard performance may begin to decline when the gap between train and test hardness becomes sufficiently large.
The remainder of the paper is organized along the following research questions:
@@ 1. RQ1:
    How Can We Measure Data Hardness? Do Different Approaches Agree?
@@ 4. RQ4:
    Our six human hardness measures and one model-based measure are empirically diverse, and model performance declines on harder test data for each measure.
@@
    LMs generalize surprisingly well from easy-to-hard data, closing 70%-100% of the performance gap between unsupervised-to-hard and hard-to-hard accuracy.
@@ 2 Related Work
Curriculum Learning. Curriculum learning has historically concerned itself with model performance on hard data points Bengio et al. (2009). Previous work in this area has argued that learning from easy data first is helpful for models later learning more complex concepts and therefore performing better on hard data Xu et al. (2020). Yet, other studies also find that a hard-then-easy ordering can work well Maharana and Bansal (2022), as can non-monotonic difficulty orderings Elgaar and Amiri (2023). Whereas curriculum learning aims to improve hard test performance by optimally ordering the training data, we simply aim to investigate how well models generalize to hard data based on the hardness of the training data. Our results suggest that pretrained LMs generalize surprisingly well from easy to hard data, potentially alleviating the need for heavily engineered training curricula.
@@
Compositional Generalization. Work in compositional generalization has previously shown that neural networks struggle to generalize to problems that require combining reasoning steps in ways not seen exactly during training Lake and Baroni (2018); Bogin et al. (2022); Zhou et al. (2023b). Further work has begun characterizing the conditions under which models generalize to compositionally more difficult problems. For instance, Transformers will generalize better on classes of algorithmic problems whose solutions can be written in RASP, meaning the programs can be implemented exactly by a Transformer forward pass Zhou et al. (2023a). Recurrent test-time computation also appears to be quite valuable for generalizing to problems requiring more reasoning steps than those seen during training Schwarzschild et al. (2021); Bansal et al. (2022). However, even GPT-3.5 with Chain-of-Thought prompting can struggle to generalize to simple mathematical problems requiring more reasoning steps than seen during training Dziri et al. (2023). Our results are not inconsistent with these studies, but instead demonstrate that relative to an unsupervised-to-hard lower bound and hard-to-hard upper bound, easy-to-hard performance on compositional reasoning problems is often surprisingly good (Sec. 5.2). In Sec. 5.1, we demonstrate the more basic conclusion (agreeing with past studies) that models perform worse on test data requiring a higher number of reasoning steps.
@@
Easy-to-Hard Generalization. Amodei et al. (2016) motivate the scalable oversight problem by pointing out how it could be challenging to give a proper reward signal to a model when it is difficult to assess if its outputs are correct. Assessing easy-to-hard generalization provides useful context for understanding the difficulty of the scalable oversight problem, as it tells us how we can expect models to generalize from a setting where we can properly supervise them to one where we cannot. Past work evaluates easy-to-hard generalization in NLP using model-based hardness measures Swayamdipta et al. (2020) and number of compositional reasoning steps required to solve a problem Fu et al. (2022). Swayamdipta et al. (2020) show that BERT models perform worse on commonsense reasoning tasks when trained on easy data rather than hard data according to a loss-based metric resembling minimum description length Perez et al. (2021). Fu et al. (2022) show a similar result with GPT3 models for StrategyQA and GSM8k, finding that prompting with “complex” examples does better than “simple” examples, where examples are categorized according to the number of reasoning steps in the annotated human reasoning chain for a problem. Relative to these works, we study easy-to-hard generalization with (1) a greater number of human hardness measures, including grade level, expert rating, required cognitive skills, question length, answer length, and number of reasoning steps, as well as a model-based metric, (2) multiple datasets involving science question answering, compositional reasoning, and mathematical reasoning, and (3) multiple model sizes for understanding scaling trends. In contrast to these works, we show that in a number of settings easy-to-hard generalization is comparable to hard-to-hard generalization.
@@
In concurrent work, Burns et al. (2023) present results on a related “weak-to-strong” generalization problem, with similar motivation to our work. They also present results specifically on easy-to-hard generalization for NLP tasks using a model-based hardness measure. In contrast to Burns et al. (2023), (1) we define our main performance metric (Supervised Gap Recovered) using an unsupervised model as the baseline performance rather than a weaker model as the baseline performance, which is important when an unsupervised stronger model will greatly outperform a supervised weaker model (as is observed in our experiments); (2) we use human hardness measures in addition to model-based hardness, which is important when the two may not correlate strongly (see our Fig. 3); and (3) we use publicly available datasets and open-source models rather than unidentified “NLP tasks” and API-gated models.
@@
Table 1: Hardness measures we use for each dataset. Grayed-out options are not present in the dataset annotations, and thus not used in our experiments.
@@ 3 Measuring Datapoint Hardness
Measuring easy-to-hard generalization requires drawing a distinction between easy and hard data, defined in terms of human ability to correctly label the data. There could be many ways to describe what makes problems harder, including that (1) only people with specialized training and knowledge can solve the problem; (2) it takes people longer to solve the problem; (3) people are less certain that their final solution is correct; (4) people with similar expertise naturally disagree about the solution to the problem, while agreeing that there is an objective solution; (5) experts know of a reliable method for obtaining the answer to a problem, but it is costly in terms of time and effort or possibly noisy in its outputs (like conducting scientific experiments). Each of these dimensions clearly reflects that hardness lies on a spectrum.
In this paper, we aim to capture the above properties in a number of specific measures we can obtain for each instance in our datasets, including:
@@
    Education/Grade Level: What level of education (possibly in a particular domain) would typically lead one to be able to answer the question?
@@
    Required Cognitive Skill: What cognitive skills are required to answer the question? This rating is based on Bloom’s taxonomy of cognitive skills, which are ranked in order of increasing complexity: (1) Remembering, (2) Understanding, (3) Applying, (4) Analyzing, and (5) Evaluating Bloom et al. (1956); Adams (2015).
@@
    Compositional Steps: Compositional reasoning is more difficult than executing individual reasoning “primitives,” since it requires first getting the primitives correct and then combining their solutions correctly. We consider how many individual reasoning steps are involved in answering a question (i.e., the number of subproblems whose solutions must be combined), according to human-annotated reasoning chains.
These notions of hardness are fundamentally human notions of hardness, but we can also measure a model-based metric for datapoint hardness. In this direction, we use a minimum-description-length (MDL) metric Voita and Titov (2020). In practice, MDL can be measured by computing a test datapoint’s average label probability across models of identical architecture trained on increasing quantities of training data for a task Perez et al. (2021). Intuitively, MDL captures how hard on average an in-distribution test datapoint is for a model to generalize to given some amount of training data. Ultimately, we use our MDL metric to capture how well a stronger model generalizes to data that is hard according to a weaker model, in order to simulate a setting where humans cannot label hard problems that they would like for a strong model to solve. Details for MDL computation are given in Appendix A.
@@
    Minimum Description Length: A model-based measure of datapoint hardness, given by the datapoint’s label probability averaged over different models trained with increasing amounts of training data. Datapoints with lower average label probabilities are considered harder.
In experiments, we use four datasets that contain instance-level annotations for some portion of these measures, as shown in Table 1.
@@ 2. $\bullet$
    MMLU Hendrycks et al. (2020): Domain-specific multiple-choice questions for a large number of domains. We subset to high school and college level math, physics, biology, chemistry, and computer science questions (termed MMLU-STEM-5). Here, grade level is high school (HS) vs. college. See Figure 2 (left) for an example. Random performance is 25%.
@@ 4. $\bullet$
    GSM8k Cobbe et al. (2021): U.S. grade school math word problems in direct answer format (i.e., no answer choices given). Random performance is 0%. The number of steps in the solution to the problem forms the “Num. Reasoning Steps” measure and is obtained from the human-annotated reasoning chain collected for each problem. See Figure 2 (right) for an example.
There are generally fewer hard datapoints than easy datapoints in our datasets, given the relative difficulty of collecting hard data. In MMLU-STEM-5, for example, there are 603 college level questions and 1143 high school questions. We show histograms for each hardness measure distribution in Appendix Fig. 20 (see Appendix C for further dataset information).

### v2

@@ Abstract
How can we train models to perform well on hard test data when hard training data is by definition difficult to label correctly? This question has been termed the scalable oversight problem and has drawn increasing attention as language models have continually improved. In this paper, we present the surprising conclusion that current pretrained language models often generalize relatively well from easy to hard data, even performing as well as oracle models finetuned on hard data. We demonstrate this kind of easy-to-hard generalization using simple finetuning methods like in-context learning, linear classifier heads, and QLoRA for seven different measures of datapoint hardness, including six empirically diverse human hardness measures (like grade level) and one model-based measure (loss-based). Furthermore, we show that even if one cares most about model performance on hard data, it can be better to collect easy data rather than hard data for finetuning, since hard data is generally noisier and costlier to collect. Our experiments use open models up to 70b in size and four publicly available question-answering datasets with questions ranging in difficulty from 3rd grade science questions to college level STEM questions and general-knowledge trivia. We conclude that easy-to-hard generalization in LMs is surprisingly strong for the tasks studied. [^1] Our code is publicly available at: https://github.com/allenai/easy-to-hard-generalization.
@@ 1 Introduction
Figure 1: A model prompted with easy data (e.g., 3rd Grade problems) does almost as well on a hard task (College problems) as a model prompted with hard data (the College bar). Results shown for Mixtral-8x7B with $k{=}10$ prompt examples, averaged over 5 random seeds.
It is difficult to supervise LMs (i.e., train LMs to give correct outputs) in specialized domains of human knowledge, because it is difficult to correctly label data in such domains. Labeling difficulty manifests itself in both time to annotate (and thus cost) and label noise Lease (2011); Northcutt et al. (2021). Labeling difficulty becomes severe when specific expertise is required Sambasivan et al. (2021). For example, for sufficiently specific physics problems, PhD holders and PhD students can make errors on as many as 40% of (objective) problems Rein et al. (2023). As more NLP benchmarks focus on challenging domain-specific tasks, having access to large human-labeled training corpora may become increasingly infeasible (e.g., existing benchmarks like MMLU Hendrycks et al. (2021) and GPQA Rein et al. (2023) do not come with training data). The question arises: how can we train models to solve hard problems when correctly labeling enough hard data for training is difficult? This problem is an example of the scalable oversight problem, which concerns how to give a good reward signal to a model when it is difficult to assess if its outputs are correct Amodei et al. (2016).
Figure 2: Supervising Llama-2-70b with easy data (left: high school level computer science problems; right: math problems with 2-3 reasoning steps) can enable generalization to hard data (left: a college level computer science problem; right: a math problem involving 8 reasoning steps). Prompts themselves are omitted for brevity.
In this paper, we study the problem of easy-to-hard generalization. Easy-to-hard generalization refers to model performance on hard test data when finetuned [^2] We use “finetuning” interchangeably with “training” and “fitting” to refer to fitting pretrained models to data via in-context learning (ICL), parameter efficient finetuning (QLoRA), or by training a linear classifier head. only on easy training data, defined according to some human hardness measure (like grade level). Since gathering data in domains like graduate level STEM fields is expensive and time-consuming, it would clearly be useful if we could improve model performance in these domains by only finetuning models on cleanly labeled data from simpler domains, like high school STEM questions. To assess how well current LMs generalize from easy to hard data, we fit models to easy data and test them on hard data (“easy-to-hard”), then compare them to an oracle upper bound and unsupervised lower bound. The oracle upper bound is a model that has access to labeled hard data for finetuning (“hard-to-hard”), while the unsupervised lower bound is a model that is prompted zero-shot to answer questions (“unsupervised-to-hard”). The metric we are interested in is the Supervision Gap Recovered (SGR):
@@
where Easy, Hard, and Unsupervised refer to model performance on hard test data when finetuned on easy data, hard data, or no data (zero-shot), respectively. This metric takes a value of 100% when finetuning on easy data is as effective as hard data, and it is 0% when a model finetuned on easy data is no better than prompting a model zero-shot.
Our main result is that pretrained language models generalize surprisingly well from easy to hard data, often performing almost as well as an oracle model finetuned on hard data (illustrated in Fig. 1). In experiments with ARC Clark et al. (2018), MMLU Hendrycks et al. (2021), GSM8k Cobbe et al. (2021), and StrategyQA Geva et al. (2021), we find that the Supervision Gap Recovered is usually between 70% and 100%, meaning that easy supervision is at least 70% as good as hard supervision for hard test performance (see Fig. 2 for example problems). These results are robust across (1) model family and scale (between 7b and 70b parameters), (2) six different human hardness measures and a model-based measure, (3) four datasets/tasks, and (4) several finetuning methods including in-context learning with and without chain-of-thought reasoning Brown et al. (2020); Wei et al. (2022), QLoRA Dettmers et al. (2023), and linear classifier heads Belinkov (2022). Overall, our results suggest that current LMs generalize relatively well to test data across human difficulty levels even when finetuned on data that is measurably easier than the test data. We hypothesize that this occurs because easy data elicits latent knowledge and skills from pretrained models in a hardness-invariant way.
We additionally demonstrate that easy supervision can outperform hard supervision when (1) within some data collection budget, a greater quantity of easy data can be collected than hard data, or (2) easy data can be labeled with lower error rates than hard data (Sec. 5.3). Lastly, we study how easy-to-hard generalization changes with model scale and the gap between train and test hardness (Sec. 5.4). The remainder of the paper is organized along the following research questions:
@@ 1. RQ1:
    How Can We Measure Data Hardness?
@@ 4. RQ4:
    Our six human hardness measures and one model-based measure are empirically diverse and model performance declines on harder test data for each measure.
@@
    LMs generalize surprisingly well from easy-to-hard data, closing 70%-100% of the gap between unsupervised and hard train data.
@@ 2 Related Work
Curriculum Learning. Curriculum learning has historically concerned itself with model performance on hard data points Bengio et al. (2009). Previous work in this area has argued that learning from easy data first is helpful for models later learning more complex concepts and therefore performing better on hard data Xu et al. (2020). Whereas curriculum learning aims to improve hard test performance by optimally ordering the training data, we simply aim to investigate how well models generalize to hard data based on the hardness of the training data. Our results suggest that pretrained LMs generalize surprisingly well from easy to hard data, potentially alleviating the need for heavily engineered training curricula.
@@
Compositional Generalization. Work in compositional generalization has previously shown that neural networks struggle to generalize to problems that require combining reasoning steps in ways not seen exactly during training Lake and Baroni (2018); Bogin et al. (2022); Zhou et al. (2023b). Further work has begun characterizing the conditions under which models generalize to compositionally more difficult problems. For instance, Transformers will generalize better on classes of algorithmic problems whose solutions can be written in RASP, meaning the programs can be implemented exactly by a Transformer forward pass Zhou et al. (2023a). Recurrent test-time computation also appears to be quite valuable for generalizing to problems requiring more reasoning steps than those seen during training Schwarzschild et al. (2021); Bansal et al. (2022). Interestingly, even GPT-3.5 with Chain-of-Thought prompting can struggle to generalize to simple mathematical problems requiring more reasoning steps than seen during finetuning Dziri et al. (2023). Our results are not inconsistent with these studies, but instead demonstrate that relative to an unsupervised-to-hard lower bound and hard-to-hard upper bound, easy-to-hard performance on compositional reasoning problems is often surprisingly good (Sec. 5.2).
@@
Easy-to-Hard Generalization. Amodei et al. (2016) motivate the scalable oversight problem by pointing out how it could be challenging to give a proper reward signal to a model when it is difficult to assess if its outputs are correct. Assessing easy-to-hard generalization provides useful context for understanding the difficulty of the scalable oversight problem, as it tells us how we can expect models to generalize from a setting where we can properly supervise them to one where we cannot. Past work evaluates easy-to-hard generalization in NLP using model-based hardness measures Swayamdipta et al. (2020) and number of compositional reasoning steps required to solve a problem Fu et al. (2023). Swayamdipta et al. (2020) show that BERT models perform worse on commonsense reasoning tasks when finetuned on easy data rather than hard data according to a loss-based metric resembling minimum description length Perez et al. (2021). Fu et al. (2023) show a similar result with GPT3 models for StrategyQA and GSM8k, finding that prompting with “complex” examples does better than “simple” examples, where examples are categorized according to the number of reasoning steps in the annotated human reasoning chain for a problem. Relative to these works, we study easy-to-hard generalization with (1) a greater number of human hardness measures, including grade level, expert rating, required cognitive skills, question length, answer length, and number of reasoning steps, as well as a model-based metric, (2) multiple datasets involving science question answering, compositional reasoning, and mathematical reasoning, and (3) multiple model sizes for understanding scaling trends. In contrast to these works, we show that in a number of settings easy-to-hard generalization is comparable to hard-to-hard generalization.
@@
In concurrent work, Burns et al. (2023) present results on a related “weak-to-strong” generalization problem, where a stronger model is finetuned using labels from a weaker model. They also explore easy-to-hard generalization for NLP tasks using a model-based hardness measure. In contrast to this work, (1) we define our main performance metric (Supervised Gap Recovered) using an unsupervised model as the baseline performance rather than a weaker model as the baseline performance, which is important when an unsupervised stronger model will greatly outperform a supervised weaker model (as is observed in our experiments); (2) we use human hardness measures in addition to model-based hardness, which is a more realistic and important setting when the two may not correlate strongly (see our Fig. 8); and (3) we use publicly available datasets and open-source models rather than unidentified “NLP tasks” and API-gated models.
@@
Table 1: Hardness measures we use for each dataset. Grayed-out options are not present in the dataset’s annotations, and thus not used in our experiments.
@@ 3 Measuring Datapoint Hardness
Measuring easy-to-hard generalization requires drawing a distinction between easy and hard data, defined in terms of human ability to correctly label the data. There could be many ways to describe what makes problems harder, including that (1) only people with specialized training and knowledge can solve the problem Lehman et al. (2019); (2) it takes people longer to solve the problem; (3) people are less certain that their final solution is correct; (4) people with similar expertise naturally disagree about the solution to the problem, while agreeing that there is an objective solution Dumitrache et al. (2018); Pavlick and Kwiatkowski (2019); Nie et al. (2020); (5) experts know of a reliable method for obtaining the answer to a problem, but it is costly in terms of time and effort or possibly noisy in its outputs (like conducting scientific experiments). In this paper, we aim to capture the above properties in a number of specific measures we can obtain for each instance in our datasets, including:
@@
    Education/Grade Level: What education level (possibly in a particular domain) would typically lead one to be able to answer the question?
@@
    Required Cognitive Skill: What cognitive skills are required by the question? This rating is based on Bloom’s cognitive skills taxonomy, in order of increasing complexity: (1) Remembering, (2) Understanding, (3) Applying, (4) Analyzing, and (5) Evaluating Bloom et al. (1956); Adams (2015).
@@
    Compositional Steps: Compositional reasoning is more difficult than executing individual reasoning “primitives.” We consider how many individual reasoning steps are involved in answering a question (i.e., the number of subproblems whose solutions must be combined), according to human-annotated reasoning chains.
@@
    Minimum Description Length: A model-based measure of hardness, measuring datapoint loss under a model family. Details for MDL computation are given in Appendix A.
The first six hardness measures are fundamentally human notions of hardness, but we can also measure a model-based metric for datapoint hardness. In this direction, the seventh measure is a minimum-description-length (MDL) metric Voita and Titov (2020). In practice, MDL can be measured by computing a test datapoint’s average label probability across models of identical architecture finetuned on increasing quantities of training data for a task Perez et al. (2021). Intuitively, MDL captures how hard on average an in-distribution test datapoint is for a model to generalize to given some amount of training data. Ultimately, we use our MDL metric to capture how well a stronger model generalizes to data that is hard according to a weaker model, in order to simulate a setting where humans cannot label hard problems that they would like for a strong model to solve.
In our experiments, we use four datasets that contain instance-level annotations for some portion of these measures, as shown in Table 1.
@@ 2. $\bullet$
    MMLU Hendrycks et al. (2021): Domain-specific multiple-choice questions for many domains. We subset to high school and college level math, physics, biology, chemistry, and computer science questions (MMLU-STEM-5). Grade level is high school (HS) vs. college. See Figure 2 (left) for an example. Random performance is 25%.
@@ 4. $\bullet$
    GSM8k Cobbe et al. (2021): U.S. grade school math word problems in direct answer format (i.e., no answer choices given). Random performance is 0%. The number of steps in a problem solution is the “Num. Reasoning Steps” measure and is obtained from the human-annotated reasoning chain collected for each problem. See Figure 2 (right) for an example.
There are generally fewer hard datapoints than easy datapoints in our datasets, given the relative difficulty of collecting hard data. In MMLU-STEM-5, for example, there are 603 college level questions and 1143 high school questions. We show histograms for each hardness measure distribution in Appendix Fig. 20. See Appendix C for further dataset information.